OpenStudy (anonymous):

@iGreen. A spherical melon has a diameter of 15 cm, including the rind. The edible portion has a diameter of 8 cm. What is the volume of the rind? Use 3.14 to approximate pi and round your answer to one decimal place.

OpenStudy (anonymous):

____cm^3

OpenStudy (anonymous):

Um, I guess we find the volume of both and subtract..

OpenStudy (anonymous):

\(V = \dfrac{4}{3} \pi r^3\) Plug in what we know: \(V = \dfrac{4}{3} (3.14) (7.5)^3\) Simplify

OpenStudy (anonymous):

7.5^3 = 421.87500

OpenStudy (anonymous):

1.3333 times 3.14 times 421.87500

OpenStudy (anonymous):

1766.20584

OpenStudy (anonymous):

Yes, now find the volume of the rind: \(V = \dfrac{4}{3} \pi r^3\) \(V = \dfrac{4}{3} (3.14) (4)^3\)

OpenStudy (anonymous):

1.3333 times 3.14 times 64

OpenStudy (anonymous):

267.939968

OpenStudy (anonymous):

Yes, now subtract 1766.20584 - 267.939968

OpenStudy (anonymous):

1498.265872

OpenStudy (anonymous):

so 1498.26587 is my answer

OpenStudy (anonymous):

Yes, so our answer is 1498.3

OpenStudy (anonymous):

Thanks! :) i have a few more

OpenStudy (anonymous):

No problem, and post them.